# Regret Analysis of Posterior Sampling-Based Expected Improvement for Bayesian Optimization

### v2

@@ 5 Experiments
We performed the experiments for the Matérn kernels $k_{\rm Mat}(\boldsymbol{x},\boldsymbol{x}^{\prime})=\frac{2^{1-\nu}}{\Gamma(\nu)}\left(\sqrt{2\nu}\frac{\|\boldsymbol{x}-\boldsymbol{x}^{\prime}\|}{\ell}\right)^{\nu}K_{\nu}\left(\sqrt{2\nu}\frac{\|\boldsymbol{x}-\boldsymbol{x}^{\prime}\|}{\ell}\right)$, where $\Gamma(\cdot)$ and $K_{\nu}(\cdot)$ are the gamma function and the modified Bessel function, and several benchmark functions from https://www.sfu.ca/~ssurjano/optimization.html. These experiments are shown in Appendix E. In the experiments for the Matérn kernels, we can observe a similar tendency to the experiments for the SE kernel. In the experiments for the benchmark functions, we confirmed that all BO methods perform well in most cases since we employed heuristic parameter settings for GP-UCB, IRGP-UCB, and GP-EI-$\mu^{\rm max}$. On the other hand, GP-TS, GP-PIMS, and GP-EIMS show good performance without heuristic tunings, and GP-PIMS and GP-EIMS show superior or comparable performance to GP-TS.
@@ 6 Conclusion
The regret analyses of GP-EI-based algorithms are fewer than those of theoretically established algorithms, such as GP-UCB and GP-TS. Existing analysis by Wang and de Freitas (2014) requires increasing the posterior variance in the actual algorithm, which can deteriorate the optimization performance. Therefore, we analyzed the posterior sampling-based GP-EI algorithm called GP-EIMS, which does not require rescaling the posterior variance. We derived the sublinear BCR upper bounds of the GP-EIMS algorithm in the Bayesian setting. We finally demonstrated the practical effectiveness of GP-EIMS through numerical experiments.
@@ Acknowledgments
This work was supported by JST ACT-X Grant Number (JPMJAX23CD and JPMJAX24C3), JST PRESTO Grant Number JPMJPR24J6, JST CREST (JPMJCR21D3, JPMJCR22N2), JST Moonshot R&D (JPMJMS2033-05), JSPS KAKENHI Grant Number (JP21H03498, JP23K19967, JP23K16943 and JP24K20847), and RIKEN Center for Advanced Intelligence Project.
@@
$\leq\frac{1}{T}2\beta_{T}^{1/2}\sum_{t=1}^{T}\sigma_{t}(\tilde{\boldsymbol{x}}_{T})$ $(\because\forall\boldsymbol{x}\in{\cal X},\forall t\in[T],\sigma_{t}(\boldsymbol{x})\geq\sigma_{T}(\boldsymbol{x}))$ (34)
=$\overset{\rm False}{=}O\left(\sqrt{\frac{\beta_{T}\gamma_{T}}{T}}\right),$ (35)
where $\hat{\boldsymbol{x}}_{T}=\operatorname*{arg\,max}_{\boldsymbol{x}\in{\cal X}}\mu_{T}(\boldsymbol{x})$, $\tilde{\boldsymbol{x}}_{T}\coloneqq\operatorname*{arg\,max}_{\boldsymbol{x}\in{\cal X}}\sigma_{T}(\boldsymbol{x})$. This result suggests that (if the equality shown as =$\overset{\rm False}{=}$ is true) the simple regret upper bound converges to 0 when $\gamma_{T}\beta_{T}=o(T)$ regardless of the BO algorithm. For example, $\gamma_{T}\beta_{T}=o(T)$ holds for Gaussian and Matérn-$\nu$ kernels with $2\nu>d$ in the Bayesian setting. This result is obviously strange.
@@ B.1 Auxiliary Lemmas
${\rm EI}\bigl{(}\mu_{t-1}(\boldsymbol{x}),\nu^{1/2}\sigma_{t-1}(\boldsymbol{x}),b\bigr{)}$ $\geq\nu^{1/2}\sigma_{t-1}(\boldsymbol{x})\tau\left(\frac{q-\beta_{t}^{1/2}}{\nu^{1/2}}\right)$ $(\because q-\beta_{t}^{1/2}\leq u)$ (45)
$\geq\nu^{1/2}\sigma_{t-1}(\boldsymbol{x})\left(\frac{q-\beta_{t}^{1/2}}{\nu^{1/2}}\right)$ $(\because\tau(c)=c+\tau(-c)\geq c)$ (46)
@@
$\geq\nu^{1/2}\sigma_{t-1}(\boldsymbol{x})\tau\left(\frac{\mu_{t-1}(\boldsymbol{x})-a}{\nu^{1/2}\sigma_{t-1}(\boldsymbol{x})}\right)$ $(\because a-b>0\text{ from assumption})$ (49)
$\geq\nu^{1/2}\sigma_{t-1}(\boldsymbol{x})\tau\left(\frac{-\beta_{t}^{1/2}}{\nu^{1/2}}\right)$ $(\because\mu_{t-1}(\boldsymbol{x})-a>-\beta_{t}^{1/2}\sigma_{t-1}(\boldsymbol{x})).$ (50)
@@
$\frac{\sigma_{t-1}(\boldsymbol{z})}{\sqrt{2\pi}}$ $\leq\sigma_{t-1}(\boldsymbol{x}_{t})\phi\left(\frac{\mu_{t-1}(\boldsymbol{x}_{t})-b}{\nu^{1/2}\sigma_{t-1}(\boldsymbol{x}_{t})}\right)$ $(\because\tau(c)<\phi(c)\text{ for }c<0)$ (58)
@@ B.2 Regret Bounds for GP-EI with Maximum Posterior Mean
$\leq C_{3}{\rm EI}\bigl{(}\mu_{t-1}(\boldsymbol{x}_{t}),\beta_{t}^{1/2}(\delta)\sigma_{t-1}(\boldsymbol{x}_{t}),\mu_{t-1}^{\rm max}\bigr{)},$ $(\boldsymbol{x}_{t}\text{ is the maximizer})$ (68)
@@
$\leq 2C_{3}\beta_{t}^{1/2}(\delta)\sigma_{t-1}(\boldsymbol{x}_{t}).$ $\bigl{(}\because\mu_{t-1}^{\rm max}>\mu_{t-1}(\boldsymbol{x}_{t})\bigr{)}.$ (70)
@@
[…]
@@
In the last inequality, we use $\sum_{t=1}^{T}\sigma_{t-1}^{2}(\boldsymbol{x}_{t})\leq C_{1}\gamma_{T}$ [Srinivas et al., 2010].
For the case (ii), we consider a finite set ${\cal X}_{t}$ with each dimension evenly divided by $\tau_{t}=\left\lceil bdrt^{2}\sqrt{\log(4ad/\delta)}\right\rceil$, such that $|{\cal X}_{t}|=\tau_{t}^{d}$. In addition, we denote $[\boldsymbol{x}]_{t}=\operatorname*{arg\,min}_{\boldsymbol{x}^{\prime}\in{\cal X}_{t}}\|\boldsymbol{x}-\boldsymbol{x}^{\prime}\|_{1}$. For this ${\cal X}_{t}$, using the union bound and Lemmas 5.5, 5.6, and 5.7 of [Srinivas et al., 2010], the following events simultaneously hold with probability at least $1-\delta$,
@@
since $\beta_{t}(\delta)\geq 2\log\left(\frac{(|{\cal X}_{t}|+1)\pi^{2}t^{2}}{3\delta}\right)$. Note that $+1$ in $|{\cal X}_{t}|+1$ is required to guarantee the confidence bound on $\boldsymbol{x}_{t}$. Therefore, we assume the above two events hold hereafter.
@@
As with the case (i), using Lemma B.1 at $[\boldsymbol{x}^{*}]_{t}$ with $a=f([\boldsymbol{x}^{*}]_{t})$, $b=\mu_{t-1}^{\rm max}$, and $\nu=\beta_{t}(\delta)$, we see that
@@
$\leq C_{3}{\rm EI}\bigl{(}\mu_{t-1}(\boldsymbol{x}_{t}),\beta_{t}^{1/2}(\delta)\sigma_{t-1}(\boldsymbol{x}_{t}),\mu_{t-1}^{\rm max}\bigr{)}.$ $(\boldsymbol{x}_{t}\text{ is the maximizer})$ (84)
@@
We consider a finite set ${\cal X}_{t}$ with each dimension evenly divided by $\tau_{t}=\left\lceil bdrt^{2}\bigl{(}\sqrt{\log(ad)}+\sqrt{\pi}/2\bigr{)}\right\rceil$, such that $|{\cal X}_{t}|=\tau_{t}^{d}$. In addition, we denote $[\boldsymbol{x}]_{t}=\operatorname*{arg\,min}_{\boldsymbol{x}^{\prime}\in{\cal X}_{t}}\|\boldsymbol{x}-\boldsymbol{x}^{\prime}\|_{1}$. Then, as with the proof of Theorem B.1 of [Takeno et al., 2023], we can obtain
@@
As with the case (i), using Lemma B.1 at $[\boldsymbol{x}^{*}]_{t}$ with $a=U_{t}([\boldsymbol{x}^{*}]_{t})$, $b=\mu_{t-1}^{\rm max}$, and $\nu=\beta_{t}$, we see that
@@ B.3 Regret Bounds for GP-EI with Maximum Posterior Mean among Evaluated Points
where we use the inequality $|f(\boldsymbol{x}_{1})|\leq\beta_{1}^{1/2}(\delta)\sqrt{k(\boldsymbol{x}_{1},\boldsymbol{x}_{1})}\leq\beta_{1}^{1/2}(\delta)$ for both cases (i) and (ii). Then, define an indices =${\cal M}=\{m(i)\mid\tilde{I}_{m(i)}(\boldsymbol{x}_{m(i)})>0,m(i)>1\}$ with ascending order $m(1)<\dots<m(M)$, where $M\coloneqq|{\cal M}|$. In addition, let $m(0)=m(1)-1\geq 1$. If $M=0$, we can easily show $\sum_{t=1}^{T}\tilde{I}_{t}(\boldsymbol{x}_{t})=\tilde{I}_{1}(\boldsymbol{x}_{1})\leq\beta_{1}^{1/2}(\delta)$. On the other hand, for the case of $M\geq 1$, we can obtain
@@
[…]
@@
as $f(\boldsymbol{x}_{1})\sim{\cal N}(0,k(\boldsymbol{x}_{1},\boldsymbol{x}_{1}))$. Then, define an indices =${\cal M}=\{m(i)\mid\tilde{I}_{m(i)}(\boldsymbol{x}_{m(i)})>0,m(i)>1\}$ with ascending order $m(1)<\dots<m(M)$, where $M\coloneqq|{\cal M}|$. In addition, let $m(0)=m(1)-1\geq 1$. If $M=0$, we can easily show $\sum_{t=1}^{T}\tilde{I}_{t}(\boldsymbol{x}_{t})=\tilde{I}_{1}(\boldsymbol{x}_{1})\leq\sqrt{\frac{2}{\pi}}$. Thus, we consider the case of $M\geq 1$. For any ${\cal M}$ under $M\geq 1$, we can obtain
@@
[…]
@@ Appendix C Proofs for GP-EIMS
Here, we show the proofs of the main paper’s theorems, lemmas, and corollaries.
@@
where $C_{1}\coloneqq 2/\log(1+\sigma^{-2})$ and the indicator function $\mathbbm{1}\{\eta_{t}\geq 0\}=1$ if $\eta_{t}\geq 0$, and 00 otherwise.
@@
=$\overset{(a)}{=}\sum_{t=1}^{T}\mathbb{E}_{{\cal D}_{t-1}}\left[\mathbb{E}_{t}\left[g^{*}_{t}-f(\boldsymbol{x}_{t})\right]\right]$ (148)
=$\overset{(b)}{=}\sum_{t=1}^{T}\mathbb{E}\left[g^{*}_{t}-\mu_{t-1}(\boldsymbol{x}_{t})\right].$ (149)
@@
$\leq\mathbb{E}\left[\sum_{t=1}^{T}\eta_{t}\mathbbm{1}\{\eta_{t}\geq 0\}\sigma_{t-1}(\boldsymbol{x}_{t})\right]$ $(\because\sigma_{t-1}(\boldsymbol{x}_{t})>0)$ (153)
[…]
@@
[…]
where $C_{1}\coloneqq 2/\log(1+\sigma^{-2})$ and the indicator function $\mathbbm{1}\{\eta_{t}\geq 0\}=1$ if $\eta_{t}\geq 0$, and 00 otherwise. For the third inequality, we use $\sum_{t=1}^{T}\sigma_{t-1}^{2}(\boldsymbol{x}_{t})\leq C_{1}\gamma_{T}$ [Lemma 5.2 in Srinivas et al., 2010]. ∎
@@
Fix $\beta>0$ and $U\leq\max_{\boldsymbol{x}\in{\cal X}}\bigl{\{}\mu_{t-1}(\boldsymbol{x})+\beta^{1/2}\sigma_{t-1}(\boldsymbol{x})\bigr{\}}$. Then, the following inequality holds:
@@
Due to the monotonicity of $\sigma_{t-1}(\boldsymbol{x})\geq\sigma_{t}(\boldsymbol{x})$, we see that $\sigma_{t-1}(\boldsymbol{x}_{t})\leq 1$. In addition, from Lemma D.1, $\sigma^{2}_{t-1}(\boldsymbol{z})\geq\frac{\sigma^{2}}{\sigma^{2}+t-1}$. Therefore,
@@
Assume the same condition as in Lemma 4.4. Then, by running GP-EIMS, the following inequality holds with probability at least $1-\delta$:
@@
From Lemma 4.4, we see that
@@
where $\boldsymbol{z}_{t}=\operatorname*{arg\,max}_{\boldsymbol{x}\in{\cal X}}g_{t}(\boldsymbol{x})$. Therefore, as a direct consequence of Lemma 4.2, we obtain that
@@
[…]
$\leq\sqrt{\frac{\sigma^{2}}{\sigma^{2}+t-1}}\sqrt{1-2\log(\delta)}.$ $(\because\log(2ad)\geq\log 2>0.69)$ (197)

### v3

@@ 5 Experiments
We performed the experiments for the Matérn kernels $k_{\rm Mat}(\boldsymbol{x},\boldsymbol{x}^{\prime})=\frac{2^{1-\nu}}{\Gamma(\nu)}\left(\sqrt{2\nu}\frac{\|\boldsymbol{x}-\boldsymbol{x}^{\prime}\|}{\ell}\right)^{\nu}K_{\nu}\left(\sqrt{2\nu}\frac{\|\boldsymbol{x}-\boldsymbol{x}^{\prime}\|}{\ell}\right)$, where $\Gamma(\cdot)$ and $K_{\nu}(\cdot)$ are the gamma function and the modified Bessel function, and several benchmark functions from https://www.sfu.ca/~ssurjano/optimization.html. These experiments are shown in Appendix E. In the experiments for the Matérn kernels, we observed a similar tendency to the experiments for the SE kernel. In the experiments for the benchmark functions, we confirmed that all BO methods perform well in most cases since we employed heuristic parameter settings for GP-UCB, IRGP-UCB, and GP-EI-$\mu^{\rm max}$. On the other hand, GP-TS, GP-PIMS, and GP-EIMS show good performance without heuristic tuning, and GP-PIMS and GP-EIMS exhibit superior or comparable performance to GP-TS.
@@ 6 Conclusion
The regret analyses of GP-EI-based algorithms are fewer than those of theoretically established algorithms, such as GP-UCB and GP-TS, despite the fact that GP-EI-based algorithms have been utilized in various applications. Existing analysis by Wang and de Freitas (2014) requires increasing the posterior variance in the actual algorithm, which can deteriorate the optimization performance. Therefore, we analyzed the posterior sampling-based GP-EI algorithm called GP-EIMS, which does not require rescaling the posterior variance. We derived the sublinear BCR upper bounds of the GP-EIMS algorithm in the Bayesian setting. We demonstrated that GP-EIMS achieves at least comparable performance compared with other baseline methods, especially theoretically guaranteed EI-based methods, without sacrificing the theoretical guarantee.
@@ Acknowledgments
This work was supported by JST ACT-X Grant Number (JPMJAX23CD and JPMJAX24C3), JST PRESTO Grant Number (JPMJPR24J6), JST CREST (JPMJCR21D3 and JPMJCR22N2), JST Moonshot R&D (JPMJMS2033-05), MEXT Program: Data Creation and Utilization Type Material Research and Development Project (Grant No. JPMXP1122712807), JSPS KAKENHI Grant Number (JP21H03498, JP23K19967, JP23K16943, JP24K20847, and JP25K03182), and RIKEN Center for Advanced Intelligence Project.
@@
$\leq\frac{1}{T}2\beta_{T}^{1/2}\sum_{t=1}^{T}\sigma_{t}(\tilde{\boldsymbol{x}}_{T})$ $(\text{because }\forall\boldsymbol{x}\in{\cal X},\forall t\in[T],\sigma_{t}(\boldsymbol{x})\geq\sigma_{T}(\boldsymbol{x}))$ (34)
$\overset{\rm False}{=}O\left(\sqrt{\frac{\beta_{T}\gamma_{T}}{T}}\right),$ (35)
where $\hat{\boldsymbol{x}}_{T}=\operatorname*{arg\,max}_{\boldsymbol{x}\in{\cal X}}\mu_{T}(\boldsymbol{x})$, $\tilde{\boldsymbol{x}}_{T}\coloneqq\operatorname*{arg\,max}_{\boldsymbol{x}\in{\cal X}}\sigma_{T}(\boldsymbol{x})$. This result suggests that (if the equality shown as $\overset{\rm False}{=}$ is true) the simple regret upper bound converges to 0 when $\gamma_{T}\beta_{T}=o(T)$ regardless of the BO algorithm. For example, $\gamma_{T}\beta_{T}=o(T)$ holds for SE and Matérn kernels. This result is obviously strange.
@@ B.1 Auxiliary Lemmas
${\rm EI}\bigl(\mu_{t-1}(\boldsymbol{x}),\nu^{1/2}\sigma_{t-1}(\boldsymbol{x}),b\bigr)$ $\geq\nu^{1/2}\sigma_{t-1}(\boldsymbol{x})\tau\left(\frac{q-\beta_{t}^{1/2}}{\nu^{1/2}}\right)$ $(\text{because }q-\beta_{t}^{1/2}\leq u)$ (45)
[…]
@@
$\geq\nu^{1/2}\sigma_{t-1}(\boldsymbol{x})\tau\left(\frac{\mu_{t-1}(\boldsymbol{x})-a}{\nu^{1/2}\sigma_{t-1}(\boldsymbol{x})}\right)$ $(\text{because }a-b>0\text{ from assumption})$ (49)
$\geq\nu^{1/2}\sigma_{t-1}(\boldsymbol{x})\tau\left(\frac{-\beta_{t}^{1/2}}{\nu^{1/2}}\right).$ $(\text{because }\mu_{t-1}(\boldsymbol{x})-a>-\beta_{t}^{1/2}\sigma_{t-1}(\boldsymbol{x}))$ (50)
@@
$\frac{\sigma_{t-1}(\boldsymbol{z})}{\sqrt{2\pi}}$ $\leq\sigma_{t-1}(\boldsymbol{x}_{t})\phi\left(\frac{\mu_{t-1}(\boldsymbol{x}_{t})-b}{\nu^{1/2}\sigma_{t-1}(\boldsymbol{x}_{t})}\right)$ $(\text{because }\tau(c)<\phi(c)\text{ for }c<0)$ (58)
@@ B.2 Regret Bounds for GP-EI with Maximum Posterior Mean
$\leq C_{3}{\rm EI}\bigl(\mu_{t-1}(\boldsymbol{x}_{t}),\beta_{t}^{1/2}(\delta)\sigma_{t-1}(\boldsymbol{x}_{t}),\mu_{t-1}^{\rm max}\bigr),$ $(\text{because }\boldsymbol{x}_{t}\text{ is the maximizer})$ (68)
@@
$\leq 2C_{3}\beta_{t}^{1/2}(\delta)\sigma_{t-1}(\boldsymbol{x}_{t}).$ $\bigl(\text{because }\mu_{t-1}^{\rm max}>\mu_{t-1}(\boldsymbol{x}_{t})\bigr).$ (70)
@@
$\leq\left(2C_{3}+\sqrt{\log(T-1+\sigma^{2})-\log(\sigma^{2})}+1\right)\sqrt{\sum_{t=1}^{T}\beta_{t}(\delta)\sum_{t=1}^{T}\sigma_{t-1}^{2}(\boldsymbol{x}_{t})}$ (76)
@@
In the second last and last inequalities, we use Cauchy–Schwarz inequality and the inequality $\sum_{t=1}^{T}\sigma_{t-1}^{2}(\boldsymbol{x}_{t})\leq C_{1}\gamma_{T}$ [Srinivas et al., 2010].
For the case (ii), we consider a finite set ${\cal X}_{t}$ with each dimension evenly divided by $\tau_{t}=\left\lceil bdrt^{2}\sqrt{\log(4ad/\delta)}\right\rceil$, such that $|{\cal X}_{t}|=\tau_{t}^{d}$. In addition, we denote $\boldsymbol{x}]_{t}=\operatorname*{arg\,min}_{\boldsymbol{x}^{\prime}\in{\cal X}_{t}}\|\boldsymbol{x}-\boldsymbol{x}^{\prime}\|_{1}$. For this ${\cal X}_{t}$, using the union bound and Lemmas 5.5, 5.6, and 5.7 of [Srinivas et al., 2010], the following events simultaneously hold with probability at least $1-\delta$,
@@
since $\beta_{t}(\delta)\geq 2\log\left(\frac{(|{\cal X}_{t}|+1)\pi^{2}t^{2}}{3\delta}\right)$. Note that +1+1 in $|{\cal X}_{t}|+1$ is required to guarantee the confidence bound on $\boldsymbol{x}_{t}$. Therefore, we assume the above two events hold hereafter.
@@
As with the case (i), using Lemma B.1 at $\boldsymbol{x}^{*}]_{t}$ with $a=f([\boldsymbol{x}^{*}]_{t})$, $b=\mu_{t-1}^{\rm max}$, and $\nu=\beta_{t}(\delta)$, we see that
@@
$\leq C_{3}{\rm EI}\bigl(\mu_{t-1}(\boldsymbol{x}_{t}),\beta_{t}^{1/2}(\delta)\sigma_{t-1}(\boldsymbol{x}_{t}),\mu_{t-1}^{\rm max}\bigr).$ $(\text{because }\boldsymbol{x}_{t}\text{ is the maximizer})$ (84)
@@
We consider a finite set ${\cal X}_{t}$ with each dimension evenly divided by $\tau_{t}=\left\lceil bdrt^{2}\bigl(\sqrt{\log(ad)}+\sqrt{\pi}/2\bigr)\right\rceil$, such that $|{\cal X}_{t}|=\tau_{t}^{d}$. In addition, we denote $\boldsymbol{x}]_{t}=\operatorname*{arg\,min}_{\boldsymbol{x}^{\prime}\in{\cal X}_{t}}\|\boldsymbol{x}-\boldsymbol{x}^{\prime}\|_{1}$. Then, as with the proof of Theorem B.1 of [Takeno et al., 2023], we can obtain
@@
As with the case (i), using Lemma B.1 at $\boldsymbol{x}^{*}]_{t}$ with $a=U_{t}([\boldsymbol{x}^{*}]_{t})$, $b=\mu_{t-1}^{\rm max}$, and $\nu=\beta_{t}$, we see that
@@ B.3 Regret Bounds for GP-EI with Maximum Posterior Mean among Evaluated Points
where we use the inequality $|f(\boldsymbol{x}_{1})|\leq\beta_{1}^{1/2}(\delta)\sqrt{k(\boldsymbol{x}_{1},\boldsymbol{x}_{1})}\leq\beta_{1}^{1/2}(\delta)$ for both cases (i) and (ii). Then, define an indices ${\cal M}=\{m(i)\mid\tilde{I}_{m(i)}(\boldsymbol{x}_{m(i)})>0,m(i)>1\}$ with ascending order $m(1)<\dots<m(M)$, where $M\coloneqq|{\cal M}|$. In addition, let $m(0)=m(1)-1\geq 1$. If $M=0$, we can easily show $\sum_{t=1}^{T}\tilde{I}_{t}(\boldsymbol{x}_{t})=\tilde{I}_{1}(\boldsymbol{x}_{1})\leq\beta_{1}^{1/2}(\delta)$. On the other hand, for the case of $M\geq 1$, we can obtain
@@
$\leq\sum_{i=1}^{M}f(\boldsymbol{x}_{m(i)})-\mu_{m(i)-1}(\boldsymbol{x}_{m(i-1)})$ $(\text{because }\text{ Definition of }\tilde{\mu}_{t-1}^{\rm max})$ (123)
@@
[…]
@@
$\mathbb{E}\left[\max_{\boldsymbol{x}\in{\cal X}}f(\boldsymbol{x})-\min_{\boldsymbol{x}\in{\cal X}}f(\boldsymbol{x})\right]$ $\leq 2\sqrt{2d\log\left(\left\lceil bdr(\log(ad)+\sqrt{\pi}/2)\right\rceil\right)+2-2\log 2}+2,$ $\text{for the case (ii)}.$ (138)
@@ Appendix C Proofs for GP-EIMS
Here, we show the proofs of the main paper’s theorems, lemmas, and corollaries. Our proof flow is summarized as follows. First, from Lemma 4.1, we can see that
${\rm BCR}_{T}$ $\leq\sqrt{\sum_{t=1}^{T}\mathbb{E}\left[\eta_{t}^{2}\mathbbm{1}\{\eta_{t}\geq 0\}\right]}\sqrt{C_{1}\gamma_{T}}.$
Then, for the case of $|{\cal X}|<\infty$, we show the upper bound of $\mathbb{E}\left[\eta_{t}^{2}\mathbbm{1}\{\eta_{t}\geq 0\}\right]=O(\log t)$ in Corollary 4.5, which is shown by Lemmas 4.2 and 4.4. Therefore, combining Lemma 4.1 and Corollary 4.5, we can obtain the sublinear BCR upper bound of $O(\sqrt{T\gamma_{T}\log T})$ as in Theorem 4.6. For the case that Assumption 2.1 holds, by adapting the credible interval obtained in Lemma 4.7, a similar proof derives the BCR upper bound of $O(\sqrt{T\gamma_{T}\log T})$ as in Theorem 4.8.
@@
where $C_{1}\coloneqq 2/\log(1+\sigma^{-2})$ and the indicator function $\mathbbm{1}\{\eta_{t}\geq 0\}=1$ if $\eta_{t}\geq 0$, and 0 otherwise.
@@
$\overset{(a)}{=}\sum_{t=1}^{T}\mathbb{E}_{{\cal D}_{t-1}}\left[\mathbb{E}_{t}\left[g^{*}_{t}-f(\boldsymbol{x}_{t})\right]\right]$ (148)
$\overset{(b)}{=}\sum_{t=1}^{T}\mathbb{E}\left[g^{*}_{t}-\mu_{t-1}(\boldsymbol{x}_{t})\right].$ (149)
@@
$\leq\mathbb{E}\left[\sum_{t=1}^{T}\eta_{t}\mathbbm{1}\{\eta_{t}\geq 0\}\sigma_{t-1}(\boldsymbol{x}_{t})\right]$ $(\text{because }\sigma_{t-1}(\boldsymbol{x}_{t})>0)$ (153)
$\leq\mathbb{E}\left[\sqrt{\sum_{t=1}^{T}\eta_{t}^{2}\mathbbm{1}\{\eta_{t}\geq 0\}\sum_{t=1}^{T}\sigma_{t-1}^{2}(\boldsymbol{x}_{t})}\right]$ $(\text{because of }\text{Cauchy--Schwarz inequality })$ (154)
@@
$\leq\sqrt{\mathbb{E}\left[\sum_{t=1}^{T}\eta_{t}^{2}\mathbbm{1}\{\eta_{t}\geq 0\}\right]}\sqrt{C_{1}\gamma_{T}},$ $\left(\text{because of }\text{Jensen's inequality}\right)$ (156)
where $C_{1}\coloneqq 2/\log(1+\sigma^{-2})$ and the indicator function $\mathbbm{1}\{\eta_{t}\geq 0\}=1$ if $\eta_{t}\geq 0$, and 0 otherwise. For the third inequality, we use $\sum_{t=1}^{T}\sigma_{t-1}^{2}(\boldsymbol{x}_{t})\leq C_{1}\gamma_{T}$ [Lemma 5.2 in Srinivas et al., 2010]. ∎
@@
Fix $\beta>0$ and $U\leq\max_{\boldsymbol{x}\in{\cal X}}\bigl\{\mu_{t-1}(\boldsymbol{x})+\beta^{1/2}\sigma_{t-1}(\boldsymbol{x})\bigr\}$. Suppose that $k(\boldsymbol{x},\boldsymbol{x})\leq 1$ for all $\boldsymbol{x}\in{\cal X}$. Then, the following inequality holds:
@@
Due to the monotonicity of $\sigma_{t-1}(\boldsymbol{x})\geq\sigma_{t}(\boldsymbol{x})$ and the assumption of $k(\boldsymbol{x},\boldsymbol{x})\leq 1$, we see that $\sigma_{t-1}(\boldsymbol{x}_{t})\leq 1$. In addition, from Lemma D.1, $\sigma^{2}_{t-1}(\boldsymbol{z})\geq\frac{\sigma^{2}}{\sigma^{2}+t-1}$. Therefore,
@@
Assume the same condition as in Lemma 4.4. Suppose that $k(\boldsymbol{x},\boldsymbol{x})\leq 1$ for all $\boldsymbol{x}\in{\cal X}$. Then, by running GP-EIMS, the following inequality holds with probability at least $1-\delta$:
@@
From Lemma 4.4, defining $\boldsymbol{z}_{t}=\operatorname*{arg\,max}_{\boldsymbol{x}\in{\cal X}}g_{t}(\boldsymbol{x})$, we see that
@@
since $g^{*}_{t}=g_{t}(\boldsymbol{z}_{t})$. Therefore, as a direct consequence of Lemma 4.2, we obtain that
@@
$\leq\sqrt{\frac{\sigma^{2}}{\sigma^{2}+t-1}}\sqrt{1-\frac{\log(\delta)}{\log(2ad)}}$ $(\text{because }\text{$m_{t}\geq bdr\sqrt{(\sigma^{2}+t-1)\log(2ad)/\sigma^{2}}$})$ (196)
$\leq\sqrt{\frac{\sigma^{2}}{\sigma^{2}+t-1}}\sqrt{1-2\log(\delta)}.$ $(\text{because }\log(2ad)\geq\log 2>0.69)$ (197)